OpenStudy (anonymous):

Jodi went 54 miles on 3.5 gallons of gas, and she wants to know how far she can go on 9 gallons of gas. Which proportion would be useful to solve the problem? A) 9/3 = 54/x B) 9/54 = x/3 C) 54/3.5 = x/9 D) 54/9 = 3.5/x

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Im pretty sure the answer is C. (:

jimthompson5910 (jim_thompson5910):

(# of miles)/(# of gallons) = (# of miles)/(# of gallons) (54 miles)/(3.5 gallons) = (x miles)/(9 gallons) 54/3.5 = x/9

OpenStudy (anonymous):

thank you!!!

jimthompson5910 (jim_thompson5910):

yw